# Graph Foundation Models

### v2

@@ Abstract
Graph Foundation Model (GFM) is a new trending research topic in the graph domain, aiming to develop a graph model capable of generalizing across different graphs and tasks. However, a versatile GFM has not yet been achieved. The key challenge in building GFM is how to enable positive transfer across graphs with diverse structural patterns. Inspired by the existing foundation models in the CV and NLP domains, we propose a novel perspective for the GFM development by advocating for a “graph vocabulary”, in which the basic transferable units underlying graphs encode the invariance on graphs. We ground the graph vocabulary construction from essential aspects including network analysis, theoretical foundations, and stability. Such a vocabulary perspective can potentially advance the future GFM design following the neural scaling laws.
@@ 1 Introduction
Foundation models (Bommasani et al., 2021), which are pre-trained on massive data and can be adapted to tackle a wide range of downstream tasks, have achieved inimitable success in various domains, e.g., computer vision (CV) (Radford et al., 2021) and natural language processing (NLP) (Bubeck et al., 2023; Touvron et al., 2023). Typically, foundation models can effectively utilize both the prior knowledge obtained from the pre-training stage and the data from downstream tasks to achieve better performance (Han et al., 2021) and even deliver promising efficacy with few-shot task demonstrations (Dong et al., 2022).
Meanwhile, graphs are vital and distinctive data structures that encapsulate non-Euclidean and intricate object relationships. Given that various graphs embody unique relations, most graph learning approaches are tailored to train from scratch for a single task on a particular graph. Such an approach requires data collection and deployment for each individual graph and task. Consequently, an intriguing question emerges: Is it possible to devise a Graph Foundation Model (GFM) capable of generalizing across different graphs and tasks?
Despite the significant advancements in foundation models in other domains, the development of GFMs remains in the infant stage. Recent research has demonstrated the potential of GFMs in specialized areas, such as knowledge graphs (Galkin et al., 2023) and molecular structures (Beaini et al., 2023). However, these models often concentrate on limited scenarios: they may be applied across multiple datasets but restricted to a single task, span various tasks within a single domain, or be confined to a small set of datasets and tasks. To date, a versatile GFM that can generalize across the broad spectrum of graph-based applications has yet to be realized.
The key difficulty in achieving the GFM is how to achieve positive transfer across diverse graph data ranging from social networks to molecule graphs with countless structure patterns. The answer from the CV and NLP domains is a shared vocabulary. In the NLP foundation models, the text is first broken down into smaller units based on the vocabulary, which can be words, phrases, or symbols. In the CV foundation models, the image is mapped to a series of discrete image tokens (Yu et al., 2023a; Bai et al., 2023) based on the vision token vocabulary. The vocabulary defines the basic units in the particular domain, transferable across different tasks and datasets. Therefore, the key challenges in achieving the GFM narrow down to how we can find the graph vocabulary, the basic transferable units underlying graphs to encode the invariance on graphs.
@@
Our contributions: In this paper, we present a vocabulary perspective to clearly state the position of the GFM. In particular, we attribute the existing success of primitive GFMs to the suitable vocabulary construction guided by the particular transferability principle on graphs in Section 2. The comprehensive review of the transferability principles on graphs is illustrated in Section 3, serving us the principle for future vocabulary construction and the GFM design.
In Section 4, we discuss the potential for building the GFM following neural scaling laws from the vocabulary perspective and review the current progress towards advanced GFM. Finally, we introduce more insights and open questions to inspire constructive discussions on the GFM in Section 5.
Comparison with past relevant literature Concurrent to our position paper, Jin et al. (2023a); Li et al. (2023b); Zhang et al. (2023c) reviews those methods adapting large language model (LLM) for graph, which haven’t shown transferring capabilities and thus diverge from our scope to build a graph-centric GFM. Liu et al. (2023e) further discusses existing graph pre-training and adaption techniques with a focus on their implementations. Instead of technical details, our work focuses more on the fundamental principles, e.g., geometric invariance across datasets. With principle guidance, we depict the promising and relatively elusive directions for the development of GFMs.
2 Existing primitive GFMs and the key reasons for the existing success
The GFM is formally defined in Liu et al. (2023e), which expects a GFM with the capability to transfer across all the graph tasks and datasets.
Nonetheless, there remains no such GFM satisfying the above definition yet. In contrast, we have observed that existing models (Galkin et al., 2023; Zheng et al., 2023a) have achieved the initial success, obtaining the desired capability of the GFM on certain aspects.
In this work, we name them as the “primitive GFM”, which can further be categorized into task-specific, domain-specific, and prototype GFMs based on their transferability. Details can be found in Section 2.1. The key reasons for their success is then further discussed in Section 2.2.
2.1 Existing primitive GFMs
@@
Based on the model transferability across domains and tasks, we can roughly distinguish the existing primitive GFMs into three categories: task-specific, domain-specific, and prototype GFMs. We provide definitions and examples for each category, with additional illustrations in Appendix B.
@@
A task-specific/domain-specific GFM should be transferable across the specific task/domain and thus adapt to diverse downstream datasets and domain-specific tasks. A notable example of a task-specific GFM is ULTRA (Galkin et al., 2023), achieving superior zero-shot knowledge graph completion (KGC) performance across datasets from various domains. A domain-specific GFM instance, DiG (Zheng et al., 2023a), learns universal representations across various chemical tasks by leveraging domain-specific knowledge.
A Prototype GFM exhibits the capability to generalize towards a limited number of datasets and tasks. A notable example is PRODIGY (Huang et al., 2023a) which can perform few-shot in-context learning over node classification and link prediction on text-attributed graphs. However, since Huang et al. (2023a) employs a shallow GNN as its core model, the scalability and flexibility to additional datasets and tasks, given its modest model size, remains uncertain.
Concurrently, there have been efforts (Li et al., 2023b) to harness LLMs for graph-related tasks. Although LLMs excel in the NLP realm, directly simply flattening graph structures and prompting LLM yields suboptimal results, evidenced by studies on both text-attributed (Chen et al., 2023b) and non-text-attributed graphs (Wang et al., 2023a). A noteworthy approach is GraphText (Zhao et al., 2023b) employs a tree-based prompt design that retains structural semantics, delivering satisfactory few-shot in-context learning performance. However, the LLM meets fundamental issues (Saparov & He, 2022; Dziri et al., 2023) on addressing graph data. We provide a further discussion on the usage of LLM in the graph domain in Section 5.
2.2 The key reasons for the existing GFM success.
@@
Despite the empirical success achieved by the primitive GFMs, the underlying reasons remain unclear. In particular, we investigate the mechanism on the ULTRA (Galkin et al., 2023), which is a task-specific GFM focusing on the knowledge graph completion (KGC) task. THe KGC task aims to infer the missing triplet (edge), denoted as $(h,r,t)$, where $r$ is a query relationship, $h$ and $t$ are the head and tail entities, respectively. The KGC model aims to answer the query $(h,r,?)$ by predicting the tail entity $t$.
@@
However, such expressive relational vocabulary only considers the pre-defined relation types which cannot generalize to the scenario with new relation types during inference. To extend the existing relational vocabulary including new relationship type, Galkin et al. (2023) constructs a graph of relations that captures fundamental interactions independent from any graph-specific relation types, serving as the second reason for its success. The graph of relations is theoretically grounded (Gao et al., 2023) which aims to learn the double permutation-equivariant representations. Such representation is equivariant to permutations of both node entities and edge relation types. Such equivariance can be an analogy to a shared relational vocabulary. It connects the new unseen relationship types to the existing ones and maps the equivariant node pairs into the same representation despite different relation types, leading to the positive transfer.
@@
In summary, we can conclude the key for ULTRA to achieve good transferability is finding a suitable vocabulary for KGC satisfying two principles: (1) The vocabulary should not be compacted, which causes distinct node pairs to share representations, leading to potential negative transfer. (2) The vocabulary should be sufficiently inclusive to map new, unknown relationships onto the existing vocabulary, potentially enabling positive transfer.
@@
The effectiveness of finding a suitable vocabulary for building the GFM can also be found in other existing primitive GFMs with the following evidence. GraphGPT (Zhao et al., 2023c) constructs a dataset-specific vocabulary where each node corresponds to a unique node ID. Notably, GraphGPT requires specific pre-training and fine-tuning on each dataset. MoleBERT (Xia et al., 2023), the foundation model for molecule graphs, manually designs a vocabulary that transforms atom attributes into chemically meaningful codes. Similarly, OFA (Liu et al., 2023b) manually converts all the node features into meaningful text descriptions.
3 The underlying transferability principle
In the last section, we investigate the key reason for the existing success of building the GFM. The key principle for building the GFM is to construct a suitable graph vocabulary, which keeps the essential invariance across datasets and tasks. Despite the existing success,
more graph transferability principles can discover different invariances and serve as guidance for constructing new suitable graph vocabulary for future GFMs.
The following discussions are organized as follows:
@@ 3 The underlying transferability principle
We first provide a general introduction to the graph transferability principles in Section 3.1. Detailed task-specific principles on node classification, link prediction, and graph classification tasks can be found in Section 3.23.3, and 3.4, respectively. We finally discuss the principles for task transferability in Section 3.5. Notably, the following discussions majorly concentrate on the transferability of the graph structure.
3.1 Graph transferability principles
@@
In this subsection, we introduce principles that enable transferability on graphs, focusing on three key aspects: network analysis, expressiveness, and stability.
@@
Moreover, there are some other challenges that would affect the deeper GNN design to include more higher-order information. A deeper GNN may suffer from performance degradation and fail to capture higher-order structural information. There are majorly three issues: (i) The oversquashing problem (Topping et al., 2021) illustrates that the node representation is insensitive to information from important but distant nodes. (ii) The oversmoothing problem (Oono & Suzuki, 2019; Cai & Wang, 2020) illustrates that more aggregations lead to the node representations converging to a unique equilibrium, which loses the distinction between different nodes. (iii) The underreaching (Barceló et al., 2020) illustrates the failure to explore, cover, or affect all relevant nodes in the graph, leading to information loss. Various techniques are proposed to identify the root causes (Di Giovanni et al., 2023; Wu et al., 2023) and solve the expressiveness issues via new GNN (Yang et al., 2021) and graph transformer (Wu et al., 2022; Müller et al., 2023) architecture designs.
@@ 3.2 Transferability principles in node classification
While homophily predominates in network analysis, it is not a universal rule. In many real-world scenarios, “opposites attract”, resulting in networks characterized by heterophily—where nodes are more likely to link with dissimilar nodes. GNNs built with the homophily principle often struggle with heterophilious networks, except in cases of “good heterophily” (Ma et al., 2021; Luan et al., 2021), where GNNs can identify and leverage consistent patterns in connections between dissimilar nodes. However, most heterophilious networks are complex and varied, posing challenges for GNNs due to their irregular and intricate interaction patterns (Luan et al., 2023; Wang et al., 2024; Mao et al., 2023a). Consequently, GNNs’ transferability, more assured in homophilous graphs, is facing significant challenges in heterophilous ones.
@@ 3.3 Transferability principles in link prediction
To consider all the possible dependencies between node pairs, we aim for the most expressive structural representation for the link prediction. This representation should be invariant if and only if links are symmetric. Zhang & Chen (2018) achieves such structural representation by incorporating node labeling features that depend on both the source and target nodes in a link. Zhang et al. (2021) further highlights the key aspects of node labeling design, including: (1) target-nodes-distinguishing, where the source and target nodes have distinct labels compared to other nodes; and (2) permutation equivariance. Node labeling methods that fulfill these criteria, such as double radius node labeling (DRNL) and zero-one (ZO) labeling, can produce the most expressive structural representations. The expressiveness representation can find the complete set of distinct relations to differentiate all non-isomorphic node pairs, thereby mitigating the risk of negative transfer in standard GNNs. Huang et al. (2023c) extends the relational Weisfeiler-Leman framework (Barcelo et al., 2022) to link prediction and incorporate the concept of labeling tricks to multi-relational graphs.
Stability. For those equally expressive structural representations, there may still be a gap in terms of their stability. For example, empirical evidence (Zhang et al., 2021) shows that GNNs with DRNL labeling outperform those with ZO labeling. From the perspective of stability, it is crucial to maintain a bounded gap in predictions for pairs under minor perturbations. Wang et al. (2021) provides a theoretical analysis identifying key properties of stable positional encoding (GNNs should be rotation and permutation equivariant to positional encodings) that enhance generalization.
3.4 Transferability principle in graph classification

### v3

# Position: Graph Foundation Models are Already Here

@@ Abstract
Graph Foundation Models (GFMs) are emerging as a significant research topic in the graph domain, aiming to develop graph models trained on extensive and diverse data to enhance their applicability across various tasks and domains. Developing GFMs presents unique challenges over traditional Graph Neural Networks (GNNs), which are typically trained from scratch for specific tasks on particular datasets The primary challenge in constructing GFMs lies in effectively leveraging vast and diverse graph data to achieve positive transfer. Drawing inspiration from existing foundation models in the CV and NLP domains, we propose a novel perspective for the GFM development by advocating for a “graph vocabulary”, in which the basic transferable units underlying graphs encode the invariance on graphs. We ground the graph vocabulary construction from essential aspects including network analysis, expressiveness, and stability. Such a vocabulary perspective can potentially advance the future GFM design in line with the neural scaling laws. All relevant resources for GFMs design can be found at here.
@@ 1 Introduction
Foundation models (Bommasani et al., 2021), which are pre-trained on massive data and can be adapted to tackle a wide range of downstream tasks, have achieved inimitable success in various domains, e.g., computer vision (CV) (Radford et al., 2021) and natural language processing (NLP) (Bubeck et al., 2023; Touvron et al., 2023). Typically, foundation models can effectively utilize both the prior knowledge obtained from the pre-training stage and the data from downstream tasks to achieve better performance (Han et al., 2021) and even deliver promising efficacy with few-shot task demonstrations (Dong et al., 2022; Mao et al., 2024).
Meanwhile, graphs are vital and distinctive data structures that encapsulate non-Euclidean and intricate object relationships. Since various graphs embody unique relations, most graph learning approaches are tailored to train from scratch for a single task on a particular graph. This approach necessitates separate data collection and deployment for each individual graph and task. Consequently, an intriguing question emerges: Is it possible to devise a Graph Foundation Model (GFM) that benefiting from large-scale training with better generalization across different domains and tasks?
Despite advanced foundation models in other domains, the development of GFMs remains in the infant stage. Recent research has demonstrated initial successes of GFMs in specialized areas, such as knowledge graphs (Galkin et al., 2023, 2024) and molecules (Beaini et al., 2023). Notably, most of these models are built on principles specific to their domains. For instance, ULTRA for knowledge graph completion (Galkin et al., 2023) draws inspiration from double equivariance for inductive link prediction (Gao et al., 2023a). However, there is still a lack of general guidance on how to build GFMs that can effectively cater to a broad spectrum of graph-based applications.
The key difficulty in designing GFMs lies in finding the invariance across diverse graph data, ranging from social networks to molecular graphs with countless structural patterns, into the same representation space to achieve positive transfer. The answer from the CV and NLP domains is a shared vocabulary. In the NLP foundation models, the text is first broken down into smaller units based on the vocabulary, which can be words, phrases, or symbols. In the CV foundation models, the image is mapped to a series of discrete image tokens (Yu et al., 2023; Bai et al., 2023) based on the vision token vocabulary. The vocabulary defines the basic units in the particular domain, transferable across different tasks and datasets. Therefore, the key challenges in achieving the GFM narrow down to how we can find the graph vocabulary, the basic transferable units underlying graphs to encode the invariance on graphs.
@@
Our contributions: In this paper, we present a vocabulary perspective to clearly state the position of the GFM. In particular, we attribute the existing success of primitive GFMs to the suitable vocabulary construction guided by the particular transferability principle on graphs in Section 2. A comprehensive review of the graph transferability principles and corresponding actionable steps is illustrated in Section 3, serving us the principle for future vocabulary construction and the GFM design. In Section 4, we discuss the potential for building the GFM following neural scaling laws from several perspectives (1) building and training vocabulary from scratch, and (2) leveraging existing LLM. Finally, we introduce more insights and open questions to inspire constructive discussions on the GFM in Section 5.
2 Existing GFMs and Key Designs
Existing GFMs (Galkin et al., 2023; Zheng et al., 2023a) have achieved initial success, including promising zero-shot generalization to unseen graphs. Based on model transferability, current GFMs can be categorized into task-specific, domain-specific, and primitive GFMs. Definitions for all categories can be found in Section 2.1. The key to a successful GFM design is further discussed in Section 2.2. Notably, none of the current GFM have the capability to transfer across all graph tasks and datasets from all domains, despite such expectations being achieved in the NLP domain (Bubeck et al., 2023; Touvron et al., 2023) with long-term effort. GFMs remain in a nascent stage with limited development. Despite the gap compared to the success in the NLP domain, GFMs have already achieved significant improvement over existing GNNs with end-to-end training on a single dataset. However, the feasibility of general GFM remains unclear with unique graph challenges. Graphs are abstract data structures which are more diverse than natural language text and images grounded in the physical world
2.1 Existing GFM Categories
@@
Based on the model transferability across domains and tasks, we can roughly distinguish the existing primitive GFMs into three categories: task-specific, domain-specific, and primitive GFMs. We provide definitions and examples for each category, with a more comprehensive illustration in Appendix B.
@@
A task-specific/domain-specific GFM should be transferable across the specific task/domain and thus adapt to diverse downstream datasets and domain-specific tasks. A notable example of a task-specific GFM is ULTRA (Galkin et al., 2023), achieving superior zero-shot knowledge graph completion performance across datasets from various domains. A task-specific GFM shows great practical benefits, as it can be trained on data-rich domains, e.g., Wikipedia knowledge graphs, and subsequently improve effectiveness in resource-limited domains, e.g., geography knowledge graph. A domain-specific GFM instance, DiG (Zheng et al., 2023a), learns universal representations across various chemical tasks by leveraging domain-specific knowledge. The domain-specific GFM is highly efficient, as one model can serve all tasks while also delivering improved effectiveness compared to single-task models.
A primitive GFM exhibits the capability to generalize towards a limited number of datasets and tasks. A notable example is OFA (Liu et al., 2023b), which is co-trained on data ranging from citation networks and molecule graphs to knowledge graphs via a unified task formation on node, link, and graph level tasks. The OFA model can achieve comparable or even better performance over the vanilla GNNs on each task. Nonetheless, OFA requires transforming all node features into text for co-training, which may not be convenient for all types of data. This co-training paradigm may also limit its generalization to unseen tasks and domains.
2.2 The Key to A Successful GFM Design.
Despite the empirical success achieved by existing GFMs, most of them are inspired by domain/task-specific principles. In this section, we aim to illustrate the common design approach using ULTRA (Galkin et al., 2023) as a showcase.
@@
ULTRA (Galkin et al., 2023) is a task-specific GFM focusing on the knowledge graph completion (KGC) task. The KGC task aims to infer the missing triplet (edge), denoted as $(h,r,t)$, where $r$ is a query relationship, $h$ and $t$ are the head and tail entities, respectively. The KGC model aims to answer the query $(h,r,?)$ by predicting the tail entity $t$.
@@
However, such expressive relational vocabulary only considers the pre-defined relation types which cannot generalize to the scenario with new relation types during inference. To extend the existing relational vocabulary including new relationship type, Galkin et al. (2023) constructs a graph of relations that captures fundamental interactions independent from any graph-specific relation types, serving as the second reason for its success. The graph of relations is theoretically grounded (Gao et al., 2023b) which aims to learn the double permutation-equivariant representations. Such representation is equivariant to permutations of both node entities and edge relation types. Such equivariance can be an analogy to a shared relational vocabulary. It connects the new unseen relationship types to the existing ones and maps the equivariant node pairs into the same representation despite different relation types, leading to the positive transfer.
@@
In summary, we can conclude the key for ULTRA to achieve good transferability is finding a suitable vocabulary for KGC satisfying two principles: (1) The vocabulary should not be compacted, which causes distinct node pairs to share representations, leading to potential negative transfer. (2) The vocabulary should be sufficiently inclusive to map new, unknown relationships onto the existing vocabulary, potentially enabling positive transfer. Notably, the vocabulary design in GFMs does not necessarily correspond to a tokenizer or an embedding layer as in the NLP domain. Instead, it can involve a model that maps graphs from different domains into the same representation space, enabling positive transfer and serving as a prerequisite for data-scaling.
@@
The effectiveness of finding a suitable vocabulary for building the GFM can also be found in other existing primitive GFMs with the following evidence. GraphGPT (Zhao et al., 2023b) constructs a dataset-specific vocabulary where each node corresponds to a unique node ID. Notably, GraphGPT requires specific pre-training and fine-tuning on each dataset. MoleBERT (Xia et al., 2023), the foundation model for molecule graphs, manually designs a vocabulary that transforms atom attributes into chemically meaningful codes.
3 Graph Transferability Principles with Actionable Steps
In the last section, we investigate the key to building an effective GFM, which lies in constructing a suitable graph vocabulary to keep the essential invariance across datasets and tasks. Despite existing successes, more graph transferability principles, identifying different invariances, can serve as guidance for constructing new suitable graph vocabulary for future GFMs. We present a few actionable next steps inspired by these principles, highlighting their potential benefits.
@@ 3 Graph Transferability Principles with Actionable Steps
The following discussions are organized as follows: We first provide a general introduction to the graph transferability principles in Section 3.1. Detailed task-specific principles on node classification, link prediction, and graph classification tasks can be found in Section 3.23.3, and 3.4, respectively. We finally discuss the principles for task transferability in Section 3.5. Notably, the following discussions majorly concentrate on the transferability of the graph structure. The discussion about techniques for aligning the feature space can be found in Appendix C.
3.1 An overview on Graph transferability principles
@@
In this subsection, we introduce principles that enable transferability on graphs, focusing on three key aspects: network analysis, expressiveness, and stability. More discussion on other principles revolving on deeper GNNs can be found in Appendix D.
@@ 3.2 Transferability Principles in Node Classification
While homophily predominates in network analysis, it is not a universal rule. In many real-world scenarios, “opposites attract”, resulting in networks characterized by heterophily—where nodes are more likely to link with dissimilar nodes. GNNs built with the homophily principle often struggle with heterophilious networks, except in cases of “good heterophily” (Ma et al., 2021; Luan et al., 2021), where GNNs can identify and leverage consistent patterns in connections between dissimilar nodes. However, most heterophilious networks are complex and varied, posing challenges for GNNs due to their irregular and intricate interaction patterns (Luan et al., 2023; Wang et al., 2024a; Mao et al., 2023a). Consequently, GNNs’ transferability, more assured in homophilous graphs, is facing significant challenges in heterophilous ones.
@@
Actionable steps inspired by principles. (Mao et al., 2023a) illustrates the network analysis principle that a single GNN can perform well on either homophily patterns or heterophily patterns, but not both. This principle provides the actionable insight for GFM design, suggesting that the graph vocabulary for homophily patterns and heterophily patterns should be modeled separately. Consequently, the model backbone for GFMs in node classification should not rely on a single GNN, which only excels on either homophilic graphs or heterophilic graphs. A better architecture design choice could be (1) an adaptive GNN with different aggregation filters for homophilic and heterophilic graphs, or (2) a graph transformer without a fixed aggregation process.
You et al. (2023) designs a spectral regularization term inspired by the network stability to address the out-of-distribution problem. Adapting spectral regularization for GFMs could be a potential next step.
@@ 3.3 Transferability Principles in Link Prediction
To consider all the possible dependencies between node pairs, we aim for the most expressive structural representation for the link prediction. This representation should be invariant if and only if links are symmetric. Zhang & Chen (2018) achieves such structural representation by incorporating node labeling features that depend on both the source and target nodes in a link. Zhang et al. (2021) further highlights the key aspects of node labeling design, including: (1) target-nodes-distinguishing, where the source and target nodes have distinct labels compared to other nodes; and (2) permutation equivariance. Node labeling methods that fulfill these criteria, such as double radius node labeling (DRNL) and zero-one (ZO) labeling, can produce the most expressive structural representations. Many other GNNs (You et al., 2021, 2019; Wang et al., 2021) can achieve similar expressiveness, serving as the potential backbone for GFM on the link prediction task. The expressiveness representation can find the complete set of distinct relations to differentiate all non-isomorphic node pairs, thereby mitigating the risk of negative transfer in standard GNNs. Huang et al. (2023c) extends the relational Weisfeiler-Leman framework (Barcelo et al., 2022) to link prediction and incorporate the concept of labeling tricks to multi-relational graphs.
Stability. For those equally expressive structural representations, there may still be a gap in terms of their stability. For example, empirical evidence (Zhang et al., 2021) shows that GNNs with DRNL labeling outperform those with ZO labeling. From the perspective of stability, it is crucial to maintain a bounded gap in predictions for pairs under minor perturbations. Wang et al. (2021) provides a theoretical analysis identifying key properties of stable positional encoding (GNNs should be rotation and permutation equivariant to positional encodings) that enhance generalization. The stable positional encoding may be directly applied towards better GFMs.
Actionable step inspired by principles. (Mao et al., 2023b) illustrates the network analysis principle concerning the incompatibility between structural proximities and feature proximity. Node pairs with high feature proximity are likely to be with low local structural proximity and vice versa. This incompatibility leads to over-emphasis on node pairs with high structural proximity while neglecting those with high feature proximity. This principle provides actionable insight for GFM design, suggesting that the graph vocabulary for feature proximity patterns and structural proximity patterns should be modeled separately. Consequently, the model backbone for GFMs in link prediction should separately encode the pairwise structural proximity and the feature proximity.
A GNN following the expressiveness principles could include all the important structural information relevant to the link prediction (Zhang et al., 2021). Dong et al. (2024) utilizes in-context learning to effective transfer expressive GNN representations to new, unseen graphs. Satisfying performance can be found across graphs from biology, transport, web, and social domains. An actionable next step could be to better utilize expressive representations for downstream graphs from specific domains.
3.4 Transferability Principles in Graph Classification